# SuperM2M: Supervised and Mixture-to-Mixture Co-Learning for Speech Enhancement and Robust ASR

### v2

@@ 4 M2M for Speech Enhancement
Fig. 2(a) illustrates M2M training for speech enhancement. The DNN takes in far-field mixtures as input and produces an estimate $\hat{S}_{q}$ for the target speaker and an estimate $\hat{V}_{q}$ for non-target signals. Each estimate is then linearly filtered via forward convolutive prediction [68] to optimize a so-called mixture-constraint loss, which encourages the filtered estimates to add up to the close-talk mixture and each far-field mixture, thereby exploiting the weak supervision afforded by close-talk and far-field mixtures for enhancement.
This section describes the DNN setup, loss functions, FCP filtering, as well as the weaknesses of M2M, which lead to the design of SuperM2M. To avoid confusion, in Table 1 we list the major hyper-parameters we will use to describe M2M.
Table 1: Major hyper-parameters of M2M training
Symbols Description Eq. Definition
$K$, $\Delta$
Past taps & prediction delay for
reference far-field mic
(6) $\widetilde{\hat{\mathbf{S}}}_{q}(t,f)=\big{[}\hat{S}_{q}(t-K+1,f),\dots,\hat{S}_{q}(t-\Delta,f)\big{]}^{\mathsf{T}}\in{\mathbb{C}}^{K-\Delta}$
$\dot{I}$, $\dot{J}$ Past & future taps for non-reference far-field mics (7) $\overline{\hat{\mathbf{S}}}_{q}(t,f)=\big{[}\hat{S}_{q}(t-\dot{I}+1,f),\dots,\hat{S}_{q}(t+\dot{J},f)\big{]}^{\mathsf{T}}\in{\mathbb{C}}^{\dot{I}+\dot{J}}$
(7) $\acute{\hat{\mathbf{V}}}_{q}(t,f)=\big{[}\hat{V}_{q}(t-\dot{I}+1,f),\dots,\hat{V}_{q}(t+\dot{J},f)\big{]}^{\mathsf{T}}\in{\mathbb{C}}^{\dot{I}+\dot{J}}$
$\ddot{I}$, $\ddot{J}$ Past & future taps for close-talk mic (8) $\breve{\hat{\mathbf{S}}}_{q}(t,f)=\big{[}\hat{S}_{q}(t-\ddot{I}+1,f),\dots,\hat{S}_{q}(t+\ddot{J},f)\big{]}^{\mathsf{T}}\in{\mathbb{C}}^{\ddot{I}+\ddot{J}}$
(8) $\grave{\hat{\mathbf{V}}}_{q}(t,f)=\big{[}\hat{V}_{q}(t-\ddot{I}+1,f),\dots,\hat{V}_{q}(t+\ddot{J},f)\big{]}^{\mathsf{T}}\in{\mathbb{C}}^{\ddot{I}+\ddot{J}}$
$\xi$
Weight flooring factor
in FCP
(12) -
$R$
Set of future taps
$\Omega=\{0,1,\dots,R\}$ to enumerate
when estimating $\ddot{J}$ on the fly
(13) -
@@ 4.1 DNN Configurations
The DNN is trained to perform complex spectral mapping [13, 14, 15], where the real and imaginary (RI) components of far-field mixtures are stacked as input features for the DNN to predict the RI components of $\hat{S}_{q}$ and $\hat{V}_{q}$. The DNN setup is described later in Section 6.5, and the loss function next. We can optionally apply iSTFT-STFT projection to $\hat{S}_{q}$ and $\hat{V}_{q}$ before loss computation (see later Section 6.7.2 for details).
@@ 4.2 Mixture-Constraint Loss
Following UNSSOR [35], M2M [40] and the objective in (3.2), we propose the following mixture-constraint (MC) loss, which regularizes the DNN estimates $\hat{S}_{q}$ and $\hat{V}_{q}$ to have them respectively approximate $S_{q}$ and $V_{q}$, by checking whether they can be utilized to reconstruct the recorded mixtures:
$\mathcal{L}_{\text{MC}}=\mathcal{L}_{\text{MC},q}+\frac{1}{P-1}\times\sum_{p=1,p\neq q}^{P}\mathcal{L}_{\text{MC},p}+\mathcal{L}_{\text{MC},0},$ (5)
where the three terms respectively follow the ones in (3.2) and are detailed next.
@@
$=\sum\limits_{t,f}\mathcal{F}\Big{(}Y_{q}(t,f),\hat{S}_{q}(t,f)+\hat{H}_{q}(t,f)+\hat{V}_{q}(t,f)\Big{)}$
$=\sum\limits_{t,f}\mathcal{F}\Big{(}Y_{q}(t,f),\hat{S}_{q}(t,f)+\hat{\mathbf{g}}_{q}(f)^{\mathsf{H}}\widetilde{\hat{\mathbf{S}}}_{q}(t,f)+\hat{V}_{q}(t,f)\Big{)}.$ (6)
In row $2$, the DNN estimates $\hat{S}_{q}$ and $\hat{V}_{q}$ are utilized to reconstruct the mixture $Y_{q}$ via $\hat{Y}_{q}=\hat{S}_{q}+\hat{H}_{q}+\hat{V}_{q}$, with the reverberation of the target speaker, $\hat{H}_{q}$, estimated by reverberating $\hat{S}_{q}$ via $\hat{H}_{q}(t,f)=\hat{\mathbf{g}}_{q}(f)^{\mathsf{H}}\widetilde{\hat{\mathbf{S}}}_{q}(t,f)$, where $\widetilde{\hat{\mathbf{S}}}_{q}(t,f)=\big{[}\hat{S}_{q}(t-K+1,f),\dots,\hat{S}_{q}(t-\Delta,f)\big{]}^{\mathsf{T}}\in{\mathbb{C}}^{K-\Delta}$ stacks a window of past T-F units with a positive prediction delay $\Delta$, and $\hat{\mathbf{g}}_{q}(f)\in{\mathbb{C}}^{K-\Delta}$ is an estimated FCP filter to be described later in Section 4.3. $\mathcal{F}(\cdot,\cdot)$, to be described in (9), is a distance function.
Similarly, $\mathcal{L}_{\text{MC},p}$, following the second term in (3.2), is the MC loss at each non-reference far-field microphone $p$:
$\mathcal{L}{}_{\text{MC},p}$ $=\sum\limits_{t,f}\mathcal{F}\Big{(}Y_{p}(t,f),\hat{Y}_{p}(t,f)\Big{)}$
@@
$=\sum\limits_{t,f}\mathcal{F}\Big{(}Y_{p}(t,f),\hat{\mathbf{h}}_{p}(f)^{\mathsf{H}}\overline{\hat{\mathbf{S}}}_{q}(t,f)+\hat{\mathbf{r}}_{p}(f)^{\mathsf{H}}\acute{\hat{\mathbf{V}}}_{q}(t,f)\Big{)},$ (7)
where $\hat{X}_{p}(t,f)=\hat{\mathbf{h}}_{p}(f)^{\mathsf{H}}\overline{\hat{\mathbf{S}}}_{q}(t,f)$, with $\overline{\hat{\mathbf{S}}}_{q}(t,f)=\big{[}\hat{S}_{q}(t-\dot{I}+1,f),\dots,\hat{S}_{q}(t+\dot{J},f)\big{]}^{\mathsf{T}}\in{\mathbb{C}}^{\dot{I}+\dot{J}}$ and $\hat{\mathbf{h}}_{p}(f)\in{\mathbb{C}}^{\dot{I}+\dot{J}}$, and $\hat{V}_{p}(t,f)=\hat{\mathbf{r}}_{p}(f)^{\mathsf{H}}\acute{\hat{\mathbf{V}}}_{q}(t,f)$, with $\acute{\hat{\mathbf{V}}}_{q}(t,f)=\big{[}\hat{V}_{q}(t-\dot{I}+1,f),\dots,\hat{V}_{q}(t+\dot{J},f)\big{]}^{\mathsf{T}}\in{\mathbb{C}}^{\dot{I}+\dot{J}}$ and $\hat{\mathbf{r}}_{p}(f)\in{\mathbb{C}}^{\dot{I}+\dot{J}}$. $\hat{\mathbf{h}}_{p}(f)$ and $\hat{\mathbf{r}}_{p}(f)$ are both estimated FCP filters to be described in Section 4.3.
Similarly, $\mathcal{L}_{\text{MC},0}$, following the third term in (3.2) and $\mathcal{L}{}_{\text{MC},p}$ in (7), is the MC loss at the close-talk microphone:
$\mathcal{L}_{\text{MC},0}$ $=\sum\limits_{t,f}\mathcal{F}\Big{(}Y_{0}(t,f),\hat{Y}_{0}(t,f)\Big{)}$
$=\sum\limits_{t,f}\mathcal{F}\Big{(}Y_{0}(t,f),\hat{X}_{0}(t,f)+\hat{V}_{0}(t,f)\Big{)}$
$=\sum\limits_{t,f}\mathcal{F}\Big{(}Y_{0}(t,f),\hat{\mathbf{h}}_{0}(f)^{\mathsf{H}}\breve{\hat{\mathbf{S}}}_{q}(t,f)+\hat{\mathbf{r}}_{0}(f)^{\mathsf{H}}\grave{\hat{\mathbf{V}}}_{q}(t,f)\Big{)},$ (8)
where $\hat{X}_{0}(t,f)=\hat{\mathbf{h}}_{0}(f)^{\mathsf{H}}\breve{\hat{\mathbf{S}}}_{q}(t,f)$, with $\breve{\hat{\mathbf{S}}}_{q}(t,f)=\big{[}\hat{S}_{q}(t-\ddot{I}+1,f),\dots,\hat{S}_{q}(t+\ddot{J},f)\big{]}^{\mathsf{T}}\in{\mathbb{C}}^{\ddot{I}+\ddot{J}}$ and $\hat{\mathbf{h}}_{0}(f)\in{\mathbb{C}}^{\ddot{I}+\ddot{J}}$, and $\hat{V}_{0}(t,f)=\hat{\mathbf{r}}_{0}(f)^{\mathsf{H}}\grave{\hat{\mathbf{V}}}_{0}(t,f)$, with $\grave{\hat{\mathbf{V}}}_{q}(t,f)=\big{[}\hat{V}_{q}(t-\ddot{I}+1,f),\dots,\hat{V}_{q}(t+\ddot{J},f)\big{]}^{\mathsf{T}}\in{\mathbb{C}}^{\ddot{I}+\ddot{J}}$ and $\hat{\mathbf{r}}_{0}(f)\in{\mathbb{C}}^{\ddot{I}+\ddot{J}}$. $\hat{\mathbf{h}}_{0}(f)$ and $\hat{\mathbf{r}}_{0}(f)$ are estimated FCP filters to be described in Section 4.3.
@@
where $r\in\{0,1,\dots,P\}$ indexes all the microphones, ||$\cdot$|| ⋅ | computes magnitude, $\mathcal{R}(\cdot)$ and $\mathcal{I}(\cdot)$ respectively extract RI components, and the denominator in (9) balances the losses at different microphones and across training mixtures.
Notice that the DNN can use all or a subset of the far-field microphone signals as the input and for loss computation. For example, we can train a monaural enhancement model by just using the reference microphone signal as the input but computing the loss on all the microphone signals.
@@ 4.3 FCP for Filter Estimation
To compute $\mathcal{L}_{\text{MC}}$, we need to first compute the linear filters (i.e., RTFs) in (6]), (7) and (8). Following UNSSOR [35], we leverage FCP [68, 70] for filter estimation, based on the DNN estimates and observed mixtures.
Assuming that the target speaker is non-moving within each utterance, we estimate, e.g., the filter $\hat{\mathbf{h}}_{0}(f)$ in (8), by solving the following problem:
$\hat{\mathbf{h}}_{0}(f)=\underset{\mathbf{h}_{0}(f)}{\text{argmin}}\sum\limits_{t}\frac{1}{\hat{\lambda}_{0}(t,f)}\Big{|}Y_{0}(t,f)-\mathbf{h}_{0}(f)^{\mathsf{H}}\breve{\hat{\mathbf{S}}}_{q}(t,f)\Big{|}^{2},$ (11)
@@
where $r$ indexes all the microphones, $\xi$ (tuned to $10^{-2}$ in this study) floors the weighting term, and $\text{max}(\cdot)$ extracts the maximum value of a power spectrogram. Notice that we compute $\hat{\lambda}$ differently for different microphones, as the energy level of each source can be very different at close-talk and far-field microphones, and deployed microphones, even if placed close to each other, often produce very different gain levels in real-world scenarios.
4.4 On-the-fly Estimation of Future Taps $\ddot{J}$ for Close-Talk Mic
@@
The hyper-parameters for the FCP filter taps in M2M training are listed in Table 1. Their ideal values are likely different for different utterances, while it is tricky and cumbersome to tune each one of them individually for each utterance. For simplicity, in UNSSOR [35] and M2M [40], the value of each hyper-parameter is configured shared for all the training utterances.
@@
This strategy could be improved for real-recorded data, such as CHiME-4 [42]. In CHiME-4, we observe that the far-field microphones are reasonably synchronized as they are placed on, and processed by, the same device, but the close-talk microphone, placed on a different device, is not accurately synchronized with the far-field microphones and their time-misalignment can be as large as $50$ ms. On the other hand, the distance between the far-field microphone array and the close-talk microphone is unknown and can vary from utterance to utterance.
@@
In this context, for simplicity, we set, for all the training utterances, $\dot{I}=\ddot{I}$ (the past taps for non-reference far-field microphones and close-talk microphone, described in Table 1) and $\dot{J}=1$ (the future taps for non-reference far-field microphones), while we propose to, at each training step, estimate $\ddot{J}$ (the future taps for close-talk microphone) for each training utterance in the mini-batch by solving the problem below, which, following the $\mathcal{L}_{\text{MC},0}$ loss in (8), enumerates a set of future taps and finds the one that leads to the best approximation of the close-talk mixture based on very short FCP filters:
@@
$\ddot{J}$ $=\underset{Z\in\Omega}{\text{argmin}}\sum\limits_{t,f}\mathcal{F}\Big{(}Y_{0}(t,f),\hat{\mathbf{h}}_{0}(f)^{\mathsf{H}}\vec{\hat{\mathbf{S}}}_{q}(t,f)+\hat{\mathbf{r}}_{0}(f)^{\mathsf{H}}\vec{\hat{\mathbf{V}}}_{q}(t,f)\Big{)},$ (13)
@@
where $\vec{\hat{\mathbf{S}}}_{q}(t,f)=\big{[}\hat{S}_{q}(t+Z-O+1,f),\dots,\hat{S}_{q}(t+Z,f)\big{]}^{\mathsf{T}}\in{\mathbb{C}}^{O}$, $\vec{\hat{\mathbf{V}}}_{q}(t,f)=\big{[}\hat{V}_{q}(t+Z-O+1,f),\dots,\hat{V}_{q}(t+Z,f)\big{]}^{\mathsf{T}}\in{\mathbb{C}}^{O}$, $O$ is set to a small value ($3$ in this study) so that the filter is short and the amount of computation spent on solving this problem is small, $\Omega=\{0,1,\dots,R\}$ denotes a set of future taps to enumerate (with $R$ tuned to $8$ for CHiME-4 to account for potentially large errors in synchronization), $Z\in\Omega$ denotes an enumerated candidate future tap, and $\hat{\mathbf{h}}_{0}(f)$ and $\hat{\mathbf{r}}_{0}(f)$ are computed in the same way as in (11).
@@
Note that we run (13) at each training step to estimate $\ddot{J}$ for each training utterance in the mini-batch. We stop gradients for the operations in (13) in the forward pass, and no back-propagation is performed for the operations in (13). The estimated $\ddot{J}$ is then used for computing the $\mathcal{L}_{\text{MC},0}$ loss in (8).

### v3

@@ 4 M2M Review, and Adapting It for Speech Enhancement
Fig. 1(a) illustrates M2M training for speech enhancement. The DNN takes in far-field mixtures as input and produces an estimate $\hat{X}_{q}$ for the target speaker and an estimate $\hat{V}_{q}$ for non-target signals. Each estimate is then linearly filtered via forward convolutive prediction [68] to optimize a so-called mixture-constraint loss [35], which encourages the filtered estimates to add up to the close-talk mixture and each far-field mixture, thereby exploiting the weak supervision afforded by close-talk and far-field mixtures for enhancement. This section describes the DNN setup, loss functions, and FCP filtering.
@@ 4.1 DNN Configurations
The DNN is trained to perform complex spectral mapping [13, 14, 15], where the real and imaginary (RI) components of far-field mixtures are stacked as input features for the DNN to predict the RI components of $\hat{X}_{q}$ and $\hat{V}_{q}$. The DNN setup is described later in Section 7.5, and the loss function next. We can optionally apply iSTFT-STFT projection to $\hat{X}_{q}$ and $\hat{V}_{q}$ before loss computation (see later Section 7.7.2 for details).
@@ 4.2 Mixture-Constraint Loss
Following UNSSOR [35], M2M [40] and the objective in (3.2), we design the following mixture-constraint (MC) loss, which regularizes the DNN estimates $\hat{X}_{q}$ and $\hat{V}_{q}$ to have them respectively approximate $X_{q}$ and $V_{q}$, by checking whether they can be utilized to reconstruct the recorded mixtures:
$\mathcal{L}_{\text{MC}}=\mathcal{L}_{\text{MC},q}+\sum_{p=0,p\neq q}^{P}\alpha_{p}\times\mathcal{L}_{\text{MC},p},$ (4)
where the two loss terms respectively follow the ones in (3.2) and will be detailed next, and $\alpha_{p}$, a weighting term, is set to $1.0$ for the close-talk microphone and to $1/(P-1)$ for non-reference far-field microphones.
@@
$=\sum\limits_{t,f}\mathcal{F}\Big{(}Y_{q}(t,f),\hat{X}_{q}(t,f)+\hat{V}_{q}(t,f)\Big{)},$ (5)
where, in row $2$, the DNN estimates $\hat{X}_{q}$ and $\hat{V}_{q}$ are utilized to reconstruct the mixture $Y_{q}$ via $\hat{Y}_{q}=\hat{X}_{q}+\hat{V}_{q}$, and $\mathcal{F}(\cdot,\cdot)$, to be described in (7), is a distance function.
Similarly, $\mathcal{L}_{\text{MC},p}$, following the second term in (3.2), is the MC loss at each non-reference microphone $p$:
=$\mathcal{L}{}_{\text{MC},p}=\sum\limits_{t,f}\mathcal{F}\Big{(}Y_{p}(t,f),\hat{Y}_{p}(t,f)\Big{)}$
@@
$=\sum\limits_{t,f}\mathcal{F}\Big{(}Y_{p}(t,f),\hat{\mathbf{h}}_{p}(f)^{\mathsf{H}}\overline{\hat{\mathbf{X}}}_{q}(t,f;p)+\hat{\mathbf{r}}_{p}(f)^{\mathsf{H}}\overline{\hat{\mathbf{V}}}_{q}(t,f;p)\Big{)},$ (6)
where $\hat{X}_{p}(t,f)=\hat{\mathbf{h}}_{p}(f)^{\mathsf{H}}\overline{\hat{\mathbf{X}}}_{q}(t,f;p)$, with $\overline{\hat{\mathbf{X}}}_{q}(t,f;p)=\big{[}\hat{X}_{q}(t-I_{p}+1,f),\dots,\hat{X}_{q}(t+J_{p},f)\big{]}^{\mathsf{T}}\in{\mathbb{C}}^{I_{p}+J_{p}}$ and $\hat{\mathbf{h}}_{p}(f)\in{\mathbb{C}}^{I_{p}+J_{p}}$, and $\hat{V}_{p}(t,f)=\hat{\mathbf{r}}_{p}(f)^{\mathsf{H}}\overline{\hat{\mathbf{V}}}_{q}(t,f;p)$, with $\overline{\hat{\mathbf{V}}}_{q}(t,f;p)=\big{[}\hat{V}_{q}(t-I_{p}+1,f),\dots,\hat{V}_{q}(t+J_{p},f)\big{]}^{\mathsf{T}}\in{\mathbb{C}}^{I_{p}+J_{p}}$ and $\hat{\mathbf{r}}_{p}(f)\in{\mathbb{C}}^{I_{p}+J_{p}}$. $\hat{\mathbf{h}}_{p}(f)$ and $\hat{\mathbf{r}}_{p}(f)$ are both estimated filters to be described in Section 4.3.
@@
where $r\in\{0,1,\dots,P\}$ indexes all the close-talk and far-field microphones, ||$\cdot$|| ⋅ | computes magnitude, $\mathcal{R}(\cdot)$ and $\mathcal{I}(\cdot)$ respectively extract RI components, and the denominator in (7) balances the losses at different microphones and across training mixtures.
Notice that the DNN can use all or a subset of the far-field microphone signals as the input and for loss computation. For example, a monaural enhancement model can be trained by just using the reference microphone signal as the input but computing the loss on all the microphone signals.
@@ 4.3 FCP for Filter Estimation
To compute $\mathcal{L}_{\text{MC}}$, we need to first compute the linear filters (i.e., RTFs) in (6). Following UNSSOR [35], we leverage FCP [68, 70] for filter estimation, based on the DNN estimates and observed mixtures.
Assuming that the target speaker is non-moving within each utterance, we estimate, e.g., the filter $\hat{\mathbf{h}}_{p}(f)$ in (6), by solving the following problem:
$\hat{\mathbf{h}}_{p}(f)=\underset{\mathbf{h}_{p}(f)}{\text{argmin}}\sum\limits_{t}\frac{\Big{|}Y_{p}(t,f)-\mathbf{h}_{p}(f)^{\mathsf{H}}\overline{\hat{\mathbf{X}}}_{q}(t,f;p)\Big{|}^{2}}{\hat{\lambda}_{p}(t,f)},$ (9)
@@
where $r$ indexes all the microphones, $\xi$ (tuned to $10^{-2}$ in this study) floors the weighting term, and $\text{max}(\cdot)$ extracts the maximum value of a power spectrogram. We compute $\hat{\lambda}$ differently for different microphones, as the energy level of each source can be very different at close-talk and far-field microphones, and deployed microphones, even if placed close to each other, often produce very different gain levels in real-world scenarios.
5 On-the-fly Estimation of Future Taps $J_{0}$ for Close-Talk Microphone
@@
The FCP filter taps used in M2M training (i.e., $I_{p}$ and $J_{p}$ for $p\in\{0,1,\dots,P\}$ and $p\neq q$) are hyperparameters to tune. Their ideal values are likely different for different utterances, while it is tricky and cumbersome to tune each one of them individually for each utterance. For simplicity, in UNSSOR [35] and M2M [40], the value of each filter tap is configured shared for all the training utterances.
@@
This strategy should be improved when dealing with real-recorded data, such as CHiME-4 [42]. In CHiME-4, we observe that the far-field microphones are reasonably synchronized as they are placed on, and processed by, the same device, but the close-talk microphone, placed on a different device, is not accurately synchronized with the far-field microphones. Their time-misalignment can be as large as $50$ ms. On the other hand, the distance between the far-field microphone array and the close-talk microphone is unknown and can vary from utterance to utterance.
@@
In this context, for simplicity, for all the training utterances, we set the past filter taps the same for all the non-reference microphones (i.e., all the $I_{p}$ for $p\in\{0,1,\dots,P\}$ and $p\neq q$ are configured the same) and set the future tap $J_{p}$ to $1$ for all the non-reference far-field microphones, while we propose to, at each training step, estimate $J_{0}$ (i.e., the future taps for the close-talk microphone) for each training utterance in the mini-batch by solving the problem below. The problem below, following the $\mathcal{L}_{\text{MC},0}$ loss in (6), enumerates a set of future taps and finds the one that leads to the best approximation of the close-talk mixture based on very short FCP filters:
@@
$J_{0}$ $=\underset{J_{0}^{\prime}\in\Omega}{\text{argmin}}\sum\limits_{t,f}\mathcal{F}\Big{(}Y_{0}(t,f),\hat{\mathbf{h}}_{0}(f)^{\mathsf{H}}\vec{\hat{\mathbf{X}}}_{q}(t,f;0)+\hat{\mathbf{r}}_{0}(f)^{\mathsf{H}}\vec{\hat{\mathbf{V}}}_{q}(t,f;0)\Big{)},$ (11)
@@
where $\vec{\hat{\mathbf{X}}}_{q}(t,f;0)=\big{[}\hat{X}_{q}(t+J_{0}^{\prime}-K+1,f),\dots,\hat{X}_{q}(t+J_{0}^{\prime},f)\big{]}^{\mathsf{T}}\in{\mathbb{C}}^{K}$, $\vec{\hat{\mathbf{V}}}_{q}(t,f;0)=\big{[}\hat{V}_{q}(t+J_{0}^{\prime}-K+1,f),\dots,\hat{V}_{q}(t+J_{0}^{\prime},f)\big{]}^{\mathsf{T}}\in{\mathbb{C}}^{K}$, $K$ is set to a small value ($3$ in this study) so that the filter is short and the amount of computation spent on solving this problem is small, $\Omega=\{0,1,\dots,R\}$ denotes a set of future taps to enumerate (with $R$ tuned to $8$ for CHiME-4 to account for potentially large errors in synchronization), $J_{0}^{\prime}\in\Omega$ denotes an enumerated candidate future tap, and $\hat{\mathbf{h}}_{0}(f)$ and $\hat{\mathbf{r}}_{0}(f)$ are computed in the same way as in (9).
@@
Note that we run (11) at each training step to estimate $J_{0}$ for each training utterance in the mini-batch. We stop gradients for the operations in (11) in the forward pass, and hence no back-propagation is performed for the operations in (11). The estimated $J_{0}$ is then used for computing the $\mathcal{L}_{\text{MC},0}$ loss in (6).
6 SuperM2M
We first point out the weaknesses of M2M training, and then propose SuperM2M to address the weaknesses. Next, we discuss the necessity of close-talk mixtures in SuperM2M.